There is a rectangle whose perimeter is 46 meters and whose area is 112 square meters. What are it's dimensions???
Help!!! How could I answer this??

Answer by edjones(8007)

A) 2L+2W=46
B) L*W=112
.
W=112/L
2L+2(112/L)=46 Substitute 112/L for W in A).
2L+224/L=46
2L^2+224=46L Multiply both sides by L to get rid of the fraction.
2L^2-46L+224=0
2(L^2-23L+112)=0
2(L-16)(L-7)=0
L=16, L=7
actually we can make L=16 and W=7 or vice versa.
Check:
16*7=112
2*16+2*7=46
Ed